Two cars 240 km apart started at the same time and traveled toward each other at the rates of 60 kph and 45 kph. In how many hrs. will they meet?
Answer by mananth(16949)

They are moving towards each other
so add up the speeds
..
60+45 = 105 kph
..
distance apart = 240 km
..
t=d/r
t=240/105
t= 2.28 hours
they will meet after 2.28 hours
..
m.ananth@hotmail.ca